OpenStudy (anonymous):

Determine whether the following vectors form a basis of R4

OpenStudy (anonymous):

\[\left(\begin{matrix}1 \\ 1\\1\\1\end{matrix}\right), \left(\begin{matrix}1 \\ -1\\1\\-1\end{matrix}\right) \left(\begin{matrix}1 \\ 2\\4\\8\end{matrix}\right) \left(\begin{matrix}1 \\ -2\\4\\-8\end{matrix}\right)\]

OpenStudy (anonymous):

I dont know how to tell if they are linearly independent vectors..

OpenStudy (anonymous):

put them all in a matrix, column by column

OpenStudy (anonymous):

multiply by constants and set it equal to 0

OpenStudy (anonymous):